How to solve 5x+10y=80
Answer by addingup(3677)

what type of solution are you looking for? You have two variables, a dependent and an independent:
5x+10y = 80
When x = 1:
5(1)+10(7.5) = 80
When x = 2:
5(2)+10(7) = 80
When x = 3:
5(3)+10(6.5) = 80
When x = 4
5(4)+10(6) = 80
What do you notice? For every 1 increase in x, y increases 1/2
:
John